OpenStudy (anonymous):

find the distance from point a (2√3,5√2) to point b (-2√3, -√2)

OpenStudy (amistre64):

distance =\[\sqrt{(x-x_1)^2+(y-y_1)^2}\]

OpenStudy (amistre64):

(2sqrt(3),5sqrt(2)) -(-2sqrt(3),-sqrt(2)) ------------------ 4sqrt(3), 6sqrt(2) sqrt(16*3 + 36*2) = sqrt(48 + 72) = sqrt(120)

OpenStudy (amistre64):

abt: 10.954

OpenStudy (amistre64):

4*30 sooo 2sqrt(30)

OpenStudy (anonymous):

thank u